Classify the events as dependent or independent: Events A and B where P(A) = 0.5, P(B) = 0.2, and P(A and B) = 0.09

Independent or Dependent?

0.5 x 0.2=0.10 which does not equal 0.09, does this mean that the correct answer is dependent?
Answer #1

The events A and B are independent if P(A and B) = P(A) * P(B)

P(A) * P(B) = 0.5 * 0.2 = 0.10

P(A and B) = 0.09

Since P(A and B) P(A) * P(B) , A and B are not independent.

That is A and B are dependent.